# Constraining the Primoridal Black Hole Abundance with Space-Based Detectors

### v1

@@ Abstract
Overdense regions can collapse into primordial black holes (PBHs) in the early universe, which are a compelling candidate for dark matter. Current constraints leave the asteroid-mass window the only possible one for PBH to account for all the dark matter, which can only be probed indirectly by the scalar-induced gravitational waves (GWs) sourced by the curvature perturbation which forms PBH. In this work, we explore the capabilities of future space-based gravitational wave detectors, including LISA, Taiji, and TianQin, to constrain such induced GWs as well as the PBH abundance. We systematically account for the width of the primordial curvature power spectrum, and find that the asteroid-mass window can be fully probed by all three space-based interferometers. If PBHs constitute the majority of dark matter, the induced GW leaves a strong signal in the mHz band with a signal-to-noise ratio of $10^{3}\sim 10^{4}$.
@@ 1 Introduction
The nature of dark matter (DM) remains one of the central problems in modern physics. Among the viable candidates, primordial black holes (PBHs), which can form through the direct gravitational collapse of rare, highly overdense regions in the early universe [1, 2, 3, 4, 5, 6], have attracted growing attention due to their natural emergence without requiring new particles beyond the Standard Model. Unlike astrophysical black holes originated from remnants of stars, PBHs can span a wide mass range. Various experimental constraints have been reported at different mass ranges, including microlensing events [7, 8, 9, 10, 11], gravitational wave (GW) mergers [12, 13], and anomalous Supernovae [14]. While none of these observations can allow PBH to fully account for dark matter, a particularly intriguing range is the asteroid-mass window of $^{-16}1010$$^{-10}M_{\odot}$, which remains the only mass range where PBHs have not yet been observationally excluded as the dominant dark matter component, primarily due to their tiny size and the optical limit [9, 15].
A population of PBHs typically requires sizeable primordial curvature perturbations, which unavoidably sources a stochastic gravitational wave background (SGWB) at quardratic order, known as scalar-induced gravitational waves (SIGWs) [16, 17, 18, 19, 20, 21]. A rough estimate suggests that the required curvature perturbation amplitude is $\mathcal{P}_{\zeta}\sim 0.01$, leading to a current GW energy density of about $\Omega_{\mathrm{GW},0}h^{2}\sim 10^{-6}\mathcal{P}_{\zeta}^{2}\sim 10^{-10}$. A wide range of existing and upcoming gravitational wave experiments are capable of probing such a stochastic background, including pulsar timing arrays (PTAs) such as NANOGrav [22, 23], EPTA [24, 25, 26], PPTA [27, 28, 29], and CPTA [30]; space-based laser interferometers such as LISA [31, 32], Taiji [33, 34], and TianQin [35, 36]; and next-generation space-based interferometers BBO  [37, 38] and DECIGO [39, 40]. With these experiments, we can probe PBHs across different mass ranges. The horizon size determines the peak frequency of SIGWs at the time of PBH formation, which also sets the PBH mass [41]
@@
This correspondence allows detectors at different frequencies to probe different PBH mass regions, for example, planet- to stellar-mass PBHs with PTAs [41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. Notably, SIGWs associated with asteroid-mass PBHs fall within the sensitivity bands of LISA/Taiji/TianQin [41, 51, 52, 53, 54, 42, 55, 56, 36]. These missions thus offer a unique opportunity to probe the asteroid-mass PBH population, offering a direct test of this scenario in which PBHs make up a significant fraction of dark matter. A simple comparison of $\Omega_{\mathrm{GW}}\sim 10^{-10}$ with the noise curve gives a detectable frequency range of space-based inteferometers $10^{-4}\sim 10^{-1}$ Hz, which can be transferred to PBH mass range of $10^{-13}\sim 10^{-10}M_{\odot}$ by (1.1). This simple estimate leaves an undetectable gap in the asteroid-mass window $10^{-16}\sim 10^{-13}M_{\odot}$. However, once we consider the spectral shape, especially the infrared tail of the SIGW, the detectable range can fully cover the asteroid-mass window. A detailed study of the detectability is necessary, especially when the spectral shape, different noise curves of these interferometers, as well as the constraints of the other experiments are considered. This is the main task of this paper.
This paper is organized as follows. In Section 2, we review our method for estimating the PBH abundance, followed by a derivation of the SIGW energy spectrum from curvature perturbations parametrized by the primordial power spectrum in Section 3. In Section 4, we compare the design specifications and noise characteristics of three space-based interferometers, including the impact of confusion foregrounds. In Section 5, we present the resulting constraints on PBHs in the asteroid-mass window and compare them with existing observational bounds. We conclude in Section 6 by summarizing the implications of our findings for PBH dark matter.
@@ 2 PBH abundance
The PBH abundance depends on the number of overdense regions and the probability to undergo gravitational collapse. The former is determined by the statistical properties of primordial curvature perturbations, while the latter—owing to the highly nonlinear nature of the collapse—typically requires threshold values calibrated through numerical simulations, which are sensitive to the profiles of the peaks. In this section, we briefly introduce our method for calculating the abundance of primordial black holes, combining the Press–Schechter approach [57] with peak profiles in the broad-spectrum case [58, 59].
@@ 2.1 Peak Profile
with $j_{0}(z)={\rm sin}\,z/z$ the zeroth-order spherical Bessel function. Here, the probability of finding a peak with a given amplitude $\mu$ follows a specific statistical distribution, which will be discussed in detail in Section 2.3.
@@ 2.2 Threshold and PBH Mass
where $\mathscr{C}_{\ell}(r)=-(4/3)r\,\partial_{r}\zeta$ is the linearized compaction function . For simplicity, we focus on the type-I fluctuations with $d(a\,r\,e^{\zeta})/dr>0$ or $\mathscr{C}_{\ell}<4/3$. [^1] For discussions on type-II fluctuation and related PBHs, see e.g. [69, 70, 71, 72]. A PBH forms if the innermost local maximum of the compaction function, $\mathscr{C}_{m}\equiv\mathscr{C}(r_{m})$ exceeds a threshold that depends on the peak profile. Numerical simulations show that this dependence can be captured by a functional form in terms of the shape parameter $q$ [73, 74, 75]:
@@
which roughly describes the width of the innermost maximum of $\mathcal{C}(r_{m})$. In general, peaks which can collapse into PBHs are mainly concentrated near the so-called critical region, where the compaction function satisfies $\mathscr{C}_{m}-\mathscr{C}_{\rm th}\lesssim 10^{-2}$. In this regime, the PBH mass obeys a scaling law [76, 77, 78, 79, 80]
@@ 2.3 PBH abundance
Then, in critical regime, the PBH abundance at its formation is
@@ 3 SIGW
Large primordial curvature perturbations can induce second-order gravitational waves through mode coupling in the early universe (see Ref. [87] for a review). For simplicity, we focus on SIGWs produced in the absence of anisotropic stress, where the metric takes the form in the conformal Newtonian gauge [20]
@@
Here, the kernel $f$ related to the evolution of scalar perturbation can be found in Eq. (3.20) in Ref. [87]. The equation of motion (3.3) can be solved by Green function method,
@@
where the kernel is defined as $I(\eta,k,u,v)\equiv\int^{\eta}d\tilde{\eta}~G(\eta,\tilde{\eta})f(\tilde{\eta},k,u,v)$, which admits a closed-form analytical expression in the radiation-dominated era given by Eq. (22) of Ref. [88]. In this era, the Bardeen potential $\Phi$ and the comoving curvature perturbation $\zeta$ are related by $\Phi=\frac{2}{3}\zeta$ for superhorizon modes $k\mathcal{H}\ll 1$. We can write down the energy spectrum of SIGWs both at the time of generation and at the present epoch
@@ 4 Detection
In this section, we review the detectability of SGWB signals in the space-based laser interferometers, including the Laser Interferometer Space Antenna (LISA), the Taiji Program in Space and the TianQin Project. These detectors, through their unique designs and technological approaches, provide an opportunity to probe GW signals around the millihertz frequency band.
@@
  • The Laser Interferometer Space Antenna (LISA) is a European Space Agency (ESA)–led space mission, developed in collaboration with NASA and member states, to detect low-frequency gravitational waves in the band from $10^{-4}{\rm Hz}$ to $10^{-1}{\rm Hz}$ [31]. LISA will consist of three identical spacecraft flying in a triangular constellation with 2.5 million-km arm lengths in a heliocentric orbit trailing the Earth, enabling precision laser interferometry to measure perturbations of spacetime. The LISA Pathfinder technology demonstrator, launched on 3rd December 2015, successfully validated key technologies for space-based gravitational wave detection—achieving free-fall performance and interferometric precision that met and exceeded the mission’s requirements, thereby demonstrating the feasibility of the full observatory concept [90]. Following this success, LISA was formally adopted as ESA’s third large-class mission (L3) under the Cosmic Vision program in January 2024, advancing into detailed design and implementation; industrial contracts for spacecraft construction and key instrument development were signed in 2025, with a planned launch in the mid-2030s to commence science operations.
@@
  • The Taiji Program in Space is a Chinese space-based gravitational wave observatory initiative led by the Chinese Academy of Sciences to detect low-frequency gravitational waves in the band from approximately $10^{-4}{\rm Hz}$ to $10^{-1}{\rm Hz}$ using laser interferometry [33]. Taiji’s mission concept comprises three spacecraft forming an equilateral triangular constellation with arm lengths of 3 million km in a heliocentric orbit, enabling sensitivity to sources such as massive black hole mergers, extreme mass ratio inspirals and stochastic backgrounds in the mid-to-low-frequency regime. Taiji Pathfinder (also known as Taiji-1), a technology demonstration satellite, was successfully launched in August 2019. In-orbit tests have verified key technologies, including laser interferometry at picometer precision, micro-thruster control at micro-Newton resolution, and gravitational reference sensing, providing critical validation for subsequent mission stages [91]. Following this success, the Taiji program aims toward the full three-satellite constellation targeted for launch in the early 2030s to commence science operations.
@@
  • The TianQin Project is a Chinese space-based gravitational wave observatory proposed to detect low-frequency gravitational waves in the band from $10^{-3}\,\mathrm{Hz}$ to $1\,\mathrm{Hz}$ [92, 93]. TianQin’s baseline design envisions three drag-free satellites in geocentric orbits with radii of $\sqrt{3}\times 10^{5}\,\mathrm{km}$ forming an equilateral triangle whose plane is oriented nearly perpendicular to the ecliptic, with the constellation’s sensitive axis pointing toward the ultracompact binary RX J0806.3+1527 as a reference source [92, 94, 93]. To mature the required technologies, TianQin-1, a single satellite technology demonstrator launched in December 2019, has successfully tested drag-free control and inertial sensing beyond its mission requirements [95, 35]. The progress of the on-orbit validation of core subsystems demonstrates steady advancement toward the realization of TianQin’s scientific objectives.
@@
LISA $2.5\times 10^{9}$ $15\cdot\sqrt{1+\left(\frac{\rm 2mHz}{f}\right)^{4}}$ $3\cdot\sqrt{1+\left(\frac{\rm 0.4mHZ}{f}\right)^{2}}\cdot\sqrt{1+\left(\frac{f}{\rm 8mHz}\right)^{4}}$
Taiji $3\times 10^{9}$ $8\cdot\sqrt{1+\left(\frac{\rm 2mHz}{f}\right)^{4}}$ $3\cdot\sqrt{1+\left(\frac{\rm 0.4mHZ}{f}\right)^{2}}\cdot\sqrt{1+\left(\frac{f}{\rm 8mHz}\right)^{4}}$
TianQin $\sqrt{3}\times 10^{8}$ 1 $1\cdot\sqrt{1+\frac{\rm 0.1mHZ}{f}}$
Using these parameters, we compute the strain sensitivity curves for the three detectors, which are shown in Fig. 2. The overall minimum of each sensitivity curve, corresponding to the most sensitive frequency band, is primarily determined by the detector arm length. Detectors with longer arms probe lower frequencies more efficiently, as the characteristic transfer frequency $f_{*}$ scales inversely with $L$. This explains why LISA and Taiji achieve their optimal sensitivity at lower frequencies compared to TianQin.
More specifically, in the low-frequency regime the acceleration noise contribution is amplified by the factor $(2\pi f)^{-4}$, which leads to a rapid degradation of the strain sensitivity as the frequency decreases. This steep low-frequency behavior represents a generic feature of space-based interferometers and sets a fundamental limitation on their sensitivity to gravitational waves on long time scales. In practice, the low-frequency performance differs slightly among LISA, Taiji, and TianQin due to their distinct acceleration noise models. To account for potential unknown disturbances at extremely low frequencies, the acceleration noise is commonly modeled with an additional low-frequency “reddening” factor, such as $\sqrt{1+({\rm 0.1mHz}/f)}$. As discussed in Ref. [98], this precautionary treatment is motivated by the fact that the low-frequency behavior of the acceleration noise is extremely difficult to assess during ground-based testing and verification, and therefore represents a conservative design choice.
@@
Meanwhile, the sensitivity of gravitational wave detectors depend not only on instrumental noise but also on astrophysical foreground [100]. Within the Milky Way, tens of millions of compact binary systems, collectively known as galactic binaries, contribute to a significant foreground noise [101]. These galactic binaries, which include white dwarf binaries and neutron star binaries, emit stochastic gravitational wave radiation. Characterized by orbital periods shorter than a few hours, these systems radiate GWs primarily in the frequency band of $0.1-10\rm mHz$ [102, 103]. This frequency range substantially overlaps with the sensitive bands of space-based detectors. Therefore, The gravitational wave background generated by white dwarf binaries is particularly noteworthy, as it acts as a prominent foreground noise that limits our ability to detect the stochastic gravitational wave background from the early universe. After accounting for individually resolvable binary events, the residual confusion foreground can be effectively modeled as a polynomial function in logarithmic frequency space
$\log_{10}S_{{\rm con},i}(f)=\sum_{k=0}^{5}a_{k,i}\left[{\rm log}_{10}\left(\frac{f}{1\ {\rm mHz}}\right)\right]^{k}\,.$ (4.2)
where the fitting coefficients $a_{k}$ are listed in Tab. 2 [104]. This approach allows for a more accurate characterization of the foreground noise, thereby enhancing our capacity to isolate and study the underlying gravitational wave signals of interest. Note that this empirical parametrization is only valid within the frequency range of $0.1\sim 6\,\rm mHz$.
@@
where $\Omega_{\rm GW,0}(f)$ is the present-day energy density spectrum of the induced gravitational waves under consideration, $T$ denotes the total observation time, $i$ marks the different experiments, and $H_{100}=100~\mathrm{km/s/Mpc}$. Throughout this work, we adopt $T=3\,\mathrm{yr}$ and require a threshold SNR $\varrho_{i}=3$ for the detection of SGWB. This procedure leads to the power-law-integrated (PLI) sensitivity curves [106], which visualize the response of the detector to a SGWB with a power-law spectrum.
@@
Figure 3: Noise spectrum for various space-based detectors, including the instrument noise (solid black) and the confusion foreground (dashed gray). The dashed gray line shows an envelop curve of $f_{\rm PBH}=1$ obtained by shifting $\mathcal{A}_{\zeta}(k_{p})$. Colored curves indicate the marginally detectable induced gravitational wave signals with SNR $\varrho_{i}=3$ for 3 years at different peak wavenumbers: $k_{p}=10^{11}$ (blue), $10^{12}$ (orange), $10^{13}$ (green), $10^{14}$ (red), and $10^{15}$ (purple) $\mathrm{Mpc}^{-1}$. From top to bottom, the rows correspond to power spectrum widths $\Delta=0$, 0.28, 0.5, and 1, respectively. From left to right, the columns show the results for TianQin, LISA, and Taiji. The corresponding power-law-integrated sensitivity curves for SNR $\varrho_{i}=3$ and observation time $T=3\,\mathrm{yr}$ are shown in gray solid curves for comparison.

### v2

# Constraining the Primordial Black Hole Abundance with Space-Based Detectors

@@ Abstract
Overdense regions can collapse into primordial black holes (PBHs) in the early universe, which are a compelling candidate for dark matter. Current constraints leave the asteroid-mass window the only possible one for PBH to account for all the dark matter, which can only be probed indirectly by the scalar-induced gravitational waves (GWs) sourced by the curvature perturbation which forms PBH. In this work, we explore the capabilities of future space-based gravitational wave detectors, including LISA, Taiji, and TianQin, to constrain such induced GWs as well as the PBH abundance. We systematically account for the width of the primordial curvature power spectrum, and find that the asteroid-mass window can be fully probed by all three space-based interferometers. If PBHs constitute the majority of dark matter, the induced GW leaves a strong signal in the mHz band with a signal-to-noise ratio of $10^{3}$$10^{4}$.
@@ 1 Introduction
The nature of dark matter (DM) remains one of the central problems in modern physics. Among the viable candidates, primordial black holes (PBHs), which can form through the direct gravitational collapse of rare, highly overdense regions in the early universe [180, 86, 46, 120, 47, 101], have attracted growing attention due to their natural emergence without requiring new particles beyond the Standard Model. Unlike astrophysical black holes originated from remnants of stars, PBHs can span a wide mass range [84]. Various experimental constraints have been reported in different mass ranges, including microlensing events [10, 162, 135, 124, 123], gravitational wave (GW) mergers [2, 1], and anomalous supernovae [156]. While none of these observations can allow PBH to fully account for dark matter, there are no robust constraint in the asteroid-mass window of $^{-16}1010$$^{-10}M_{\odot}$. The femtolensing constraints [28, 97, 95] and dynamical constraints [43, 42, 61, 83, 122, 59] are either model-dependent or reply on future experiments (see e.g. [48] for a review). Therefore, the asteroid-mass windows remains the only mass range where PBHs have not yet been observationally excluded as the dominant dark matter component, primarily due to their tiny size [135, 158].
A non-negligible population of PBHs typically requires sizeable primordial curvature perturbation, which unavoidably sources a stochastic gravitational wave background (SGWB) at quadratic order, known as scalar-induced gravitational waves (SIGWs) [164, 117, 118, 116, 12, 31]. A rough estimate suggests that the required curvature perturbation amplitude is $\mathcal{P}_{\zeta}\sim 0.01$, leading to a current GW energy density of about $\Omega_{\mathrm{GW},0}h^{2}\sim 10^{-6}\mathcal{P}_{\zeta}^{2}\sim 10^{-10}$. A wide range of existing and upcoming gravitational wave experiments are capable of probing such a stochastic background, including pulsar timing arrays (PTAs) such as NANOGrav [5, 6], EPTA [15, 16, 17], PPTA [148, 181, 149], and CPTA [175]; space-based laser interferometers such as LISA [11, 23], Taiji [114, 173], and TianQin [113, 112]; and next-generation space-based interferometers BBO [56, 53] and DECIGO [98, 99]. With these experiments, we can probe PBHs across different mass ranges. The horizon size determines the peak frequency of SIGWs at the time of PBH formation, which also sets the PBH mass [151]
@@
This correspondence allows detectors at different frequencies to probe different PBH mass regions, for example, planet- to stellar-mass PBHs with PTAs [151, 37, 40, 51, 65, 89, 60, 63, 79, 92]. Notably, SIGWs associated with asteroid-mass PBHs fall within the sensitivity bands of LISA/Taiji/TianQin [151, 38, 167, 29, 30, 37, 150, 24, 112, 91]. These missions thus offer a unique opportunity to probe the asteroid-mass PBH population, providing a direct test of this scenario in which PBHs make up a significant fraction of dark matter. A simple comparison of $\Omega_{\mathrm{GW}}\sim 10^{-10}$ with the noise curve gives a detectable frequency range of space-based interferometers $10^{-4}$$10^{-1}$ Hz, which can be transferred to the PBH mass range of $^{-13}1010$$^{-10}M_{\odot}$ by (1.1). This simple estimate leaves an undetectable gap in the asteroid-mass window $10^{-16}$$10^{-13}M_{\odot}$. However, once we consider the spectral shape, especially the infrared tail of the SIGW, the detectable range can fully cover the asteroid-mass window. A detailed study of the detectability is necessary, especially when the spectral shape, the different noise curves of these interferometers, as well as constraints from the other experiments are considered. This is the main task of this paper.
This paper is organized as follows. In Section 2, we review our method for estimating the PBH abundance, followed by a derivation of the SIGW energy spectrum from curvature perturbations parameterized by the primordial power spectrum in Section 3. In Section 4, we compare the design specifications and noise characteristics of three space-based interferometers, including the impact of confusion foregrounds. In Section 5, we present the resulting constraints on PBHs in the asteroid-mass window and compare them with existing observational bounds. We conclude in Section 6 by summarizing the implications of our findings for PBH dark matter.
@@ 2 PBH abundance
The PBH abundance depends on the number of overdense regions and the probability of undergoing gravitational collapse. The former is determined by the statistical properties of primordial curvature perturbations, while the latter—owing to the highly nonlinear nature of the collapse—typically requires threshold values calibrated through numerical simulations, which are sensitive to the profiles of the peaks. In this section, we briefly introduce our method for calculating the abundance of primordial black holes, combining the Press–Schechter approach [179] with peak profiles in the broad-spectrum case [27, 138].
@@ 2.1 Peak Profile
with $j_{0}(z)=\sin z/z$ the zeroth-order spherical Bessel function. Here, the probability of finding a peak with a given amplitude $\mu$ follows a specific statistical distribution, which will be discussed in detail in Section 2.3.
@@ 2.2 Threshold and PBH Mass
where $\mathscr{C}_{\ell}(r)=-(4/3)r\,\partial_{r}\zeta$ is the linearized compaction function. For simplicity, we focus on the type-I fluctuations with $d(a\,r\,e^{\zeta})/dr>0$ or $\mathscr{C}_{\ell}<4/3$. [^1] For discussions on type-II fluctuation and related PBHs, see e.g. [165, 155, 166, 69, 72, 73]. A PBH forms if the innermost local maximum of the compaction function, $\mathscr{C}_{m}\equiv\mathscr{C}(r_{m})$ exceeds a threshold that depends on the peak profile. Numerical simulations show that this dependence can be captured by a functional form in terms of the shape parameter $q$ [70, 126, 81]:
@@
which roughly describes the width of the innermost maximum of $\mathscr{C}(r_{m})$. In general, peaks which can collapse into PBHs are mainly concentrated near the so-called critical region, where the compaction function satisfies $\mathscr{C}_{m}-\mathscr{C}_{\rm th}\lesssim 10^{-2}$. In this regime, the PBH mass obeys a scaling law [52, 76, 104, 134, 128]
@@ 2.3 PBH abundance
Then, in the critical regime, the PBH abundance at its formation is
@@ 3 SIGW
Large primordial curvature perturbations can induce second-order gravitational waves through mode coupling in the early universe (see Ref. [67] for a review). It has recently been shown that tensor perturbations in the Newton gauge are directly related to the time delay measured by GW detectors [64]. Throughout this work, we therefore adopt this gauge and write the perturbed metric as
@@
Here, the kernel $f$ related to the evolution of scalar perturbation can be found in Eq. (3.20) in Ref. [67]. The equation of motion (3.3) can be solved by the Green function method,
@@
where the kernel is defined as $I(\eta,k,u,v)\equiv\int^{\eta}d\tilde{\eta}~G(\eta,\tilde{\eta})f(\tilde{\eta},k,u,v)$, which admits a closed-form analytical expression in the radiation-dominated era given by Eq. (22) of Ref. [103]. (See also [74].) In this era, the Bardeen potential $\Phi$ and the comoving curvature perturbation $\zeta$ are related by $\Phi=\frac{2}{3}\zeta$ for superhorizon modes $k\mathcal{H}\ll 1$. We can write down the energy spectrum of SIGWs both at the time of generation and at the present epoch
@@ 4 Detection
In this section, we review the detectability of SGWB signals in the space-based laser interferometers, including the Laser Interferometer Space Antenna (LISA), the Taiji Program in Space, and the TianQin Project. These detectors, through their unique designs and technological approaches, provide an opportunity to probe GW signals around the millihertz frequency band.
@@
  • The Laser Interferometer Space Antenna (LISA) is a European Space Agency (ESA)–led space mission, developed in collaboration with NASA and member states, to detect low-frequency gravitational waves in the band from $10^{-4}{\rm Hz}$ to $0^{-1}{\rm Hz} [11$]. LISA will consist of three identical spacecraft flying in a triangular constellation with arm lengths of 2.5 million km in a heliocentric orbit trailing the Earth, enabling precision laser interferometry to measure perturbations of spacetime. The LISA Pathfinder technology demonstrator, launched on 3rd December 2015, successfully validated key technologies for space-based gravitational wave detection—achieving free-fall performance and interferometric precision that met and exceeded the mission’s requirements, thereby demonstrating the feasibility of the full observatory concept [18]. Following this success, LISA was formally adopted as ESA’s third large-class mission (L3) under the Cosmic Vision program in January 2024, advancing into detailed design and implementation; industrial contracts for spacecraft construction and key instrument development were signed in 2025, with a planned launch in the mid-2030s to commence science operations.
@@
  • The Taiji Program in Space is a Chinese space-based gravitational wave observatory initiative led by the Chinese Academy of Sciences to detect low-frequency gravitational waves in the band from approximately $10^{-4}{\rm Hz}$ to $10^{-1}{\rm Hz}$ using laser interferometry [114]. Taiji’s mission concept comprises three spacecraft forming an equilateral triangular constellation with arm lengths of 3 million km in a heliocentric orbit, enabling sensitivity to sources such as massive black hole mergers, extreme mass ratio inspirals, and stochastic backgrounds in the mid-to-low-frequency regime. Taiji Pathfinder (also known as Taiji-1), a technology demonstration satellite, was successfully launched in August 2019. In-orbit tests have verified key technologies, including laser interferometry at picometer precision, micro-thruster control at micro-Newton resolution, and gravitational reference sensing, providing critical validation for subsequent mission stages [174]. Following this success, the Taiji program aims toward the full three-satellite constellation targeted for launch in the early 2030s to commence science operations.
@@
  • The TianQin Project is a Chinese space-based gravitational wave observatory proposed to detect low-frequency gravitational waves in the band from $10^{-3}\,\mathrm{Hz}$ to $\mathrm{Hz} [11$0, 119]. TianQin’s baseline design envisions three drag-free satellites in geocentric orbits with radii of $\sqrt{3}\times 10^{5}\,\mathrm{km}$ forming an equilateral triangle whose plane is oriented nearly perpendicular to the ecliptic, with the constellation’s sensitive axis pointing toward the ultracompact binary RX J0806.3+1527 as a reference source [110, 109, 119]. To mature the required technologies, TianQin-1, a single satellite technology demonstrator launched in December 2019, has successfully tested drag-free control and inertial sensing beyond its mission requirements [111, 113]. The progress of the on-orbit validation of core subsystems demonstrates steady advancement toward the realization of TianQin’s scientific objectives.
@@
LISA $2.5\times 10^{9}$ $15\cdot\sqrt{1+\left(\frac{2\,\rm mHz}{f}\right)^{4}}$ $3\cdot\sqrt{1+\left(\frac{0.4\,\rm mHz}{f}\right)^{2}}\cdot\sqrt{1+\left(\frac{f}{8\,\rm mHz}\right)^{4}}$
Taiji $3\times 10^{9}$ $8\cdot\sqrt{1+\left(\frac{2\,\rm mHz}{f}\right)^{4}}$ $3\cdot\sqrt{1+\left(\frac{0.4\,\rm mHZ}{f}\right)^{2}}\cdot\sqrt{1+\left(\frac{f}{8\,\rm mHz}\right)^{4}}$
TianQin $\sqrt{3}\times 10^{8}$ 1 $1\cdot\sqrt{1+\frac{0.1\,\rm mHZ}{f}}$
Using these parameters, we compute the strain sensitivity curves for the three detectors, which are shown in Fig. 2. The overall minimum of each sensitivity curve, corresponding to the most sensitive frequency band, is primarily determined by the detector arm length. Detectors with longer arms probe lower frequencies more efficiently, as the characteristic transfer frequency $f_{*}$ scales inversely with $L$. This explains why LISA and Taiji achieve their optimal sensitivity at lower frequencies compared to TianQin. More specifically, in the low-frequency regime the acceleration noise contribution is amplified by the factor $(2\pi f)^{-4}$, which leads to a rapid degradation of the strain sensitivity as the frequency decreases. This steep low-frequency behavior represents a generic feature of space-based interferometers and sets a fundamental limitation on their sensitivity to gravitational waves on long time scales. In practice, the low-frequency performance differs slightly among LISA, Taiji, and TianQin due to their distinct acceleration noise models. To account for potential unknown disturbances at extremely low frequencies, the acceleration noise is commonly modeled with an additional low-frequency reddening factor, such as $\sqrt{1+({0.1\,\rm mHz}/f)}$. As discussed in Ref. [75], this precautionary treatment is motivated by the fact that the low-frequency behavior of the acceleration noise is extremely difficult to assess during ground-based testing and verification, and therefore represents a conservative design choice.
@@
Meanwhile, the sensitivity of gravitational wave detectors depends not only on instrumental noise but also on astrophysical foreground [96]. Within the Milky Way, tens of millions of compact binary systems, collectively known as galactic binaries, contribute to a significant foreground noise [107]. These galactic binaries, which include white dwarf binaries and neutron star binaries, emit stochastic gravitational wave radiation. Characterized by orbital periods shorter than a few hours, these systems radiate GWs primarily in the frequency band of $0.1-10\rm mHz$ [132, 133]. This frequency range substantially overlaps with the sensitive bands of space-based detectors. Therefore, the gravitational wave background generated by white dwarf binaries is particularly noteworthy, as it acts as a prominent foreground noise that limits our ability to detect the stochastic gravitational wave background from the early universe. After accounting for individually resolvable binary events, the residual confusion foreground can be effectively modeled as a polynomial function in logarithmic frequency space
$\log_{10}S_{{\rm con},i}(f)=\sum_{k=0}^{5}a_{k,i}\left[{\log}_{10}\left(\frac{f}{\rm 1\,mHz}\right)\right]^{k}\,.$ (4.2)
[…]
@@
where $\Omega_{\rm GW,0}(f)$ is the present-day energy density spectrum of the induced gravitational waves under consideration, $T$ denotes the total observation time, $i$ marks the different experiments, and $H_{100}=100\,\mathrm{km\cdot s^{-1}\cdot Mpc^{-1}}$. Throughout this work, we adopt $T=3\,\mathrm{yr}$ and require a threshold SNR $\varrho_{i}=3$ for the detection of SGWB. This procedure leads to the power-law-integrated (PLI) sensitivity curves [160], which visualize the response of the detector to a SGWB with a power-law spectrum.
@@
Figure 3: Noise spectrum for various space-based detectors, including the instrument noise (solid black) and the confusion foreground (dotted gray). The dashed gray line shows an envelope curve of $f_{\rm PBH}=1$ obtained by shifting $\mathcal{A}_{\zeta}(k_{p})$. Colored curves indicate the marginally detectable induced gravitational wave signals with SNR $\varrho_{i}=3$ for 3 years at different peak wavenumbers: $k_{p}=10^{11}$ (blue), $10^{12}$ (orange), $10^{13}$ (green), $10^{14}$ (red), and $10^{15}$ (purple) $\mathrm{Mpc}^{-1}$. From top to bottom, the rows correspond to power spectrum widths $\Delta=0$, 0.28, 0.5, and 1, respectively. From left to right, the columns show the results for TianQin, LISA, and Taiji. The corresponding power-law-integrated sensitivity curves for SNR $\varrho_{i}=3$ and observation time $T=3\,\mathrm{yr}$ are shown in gray solid curves for comparison.